OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge \frac{d}{dx} \cot^2 (\sin \theta)\]

OpenStudy (lgbasallote):

im thinking \[2\cot (\sin \theta) \cos \theta\]

OpenStudy (anonymous):

chain rule three times, just grind it out

OpenStudy (lgbasallote):

but i think there should be sec^2 somewhere...

OpenStudy (lgbasallote):

im having doubts

OpenStudy (anonymous):

you forgot to take the derivative of cotangent

OpenStudy (lgbasallote):

ohh so there really has to be a sec^2

OpenStudy (anonymous):

Yes.. -cosec^2x

OpenStudy (lgbasallote):

oops csc..what i meant haha

OpenStudy (lgbasallote):

\[-2\cot (\sin \theta\) \csc^2 (\sin \theta) \cos \theta?\]

OpenStudy (anonymous):

And that too with negative sign ha haha..

OpenStudy (lgbasallote):

yeahhh minor details lol

OpenStudy (anonymous):

Yes you are right now..

OpenStudy (lgbasallote):

nice nice

OpenStudy (anonymous):

\[-2\cot (\sin \theta). \csc^2 (\sin \theta) .\cos \theta\]